Almonds and almond extracts are common ingredients in baked goods. Almonds contain the compound benzaldehyde, shown below, which accounts for the odor of the nut. Benzaldehyde is also responsible for the aroma of cherries. Use the data provided to calculate benzaldehyde’s heat of vaporization. answer in kj/mol
Vapor Pressure (torr) Temperature (K)
(K)
56.8 376.0
602 443.0
1
Expert's answer
2020-02-07T05:28:58-0500


We solve by the following formula:


H=(R*T1*T2*lnp1/p2)/(T2-T1),


R - Universal gas constant: 8.314 j/mol*K

T2 and T1 - temperature, K

p1 and p2 - pascal pressure

p1=7572,711Pa

p2=80 260.1Pa

 

H=((8.314j/mol*K*376K*443K*(ln80,26 *103Pa/7,57*103Pa))/(443K-376K)=8.314*376*443*ln10.60/67=


8.314*376*443*2.36/67=48 779.66 j/mol =48,8 kj/mol